A car can be produced by 20 workers in the United States, and by 10 workers in Germany. 1,000 chickens can be produced by 25 workers in the United States, and by 20 workers in Germany.

a) Using charts, show how a shift of workers toward one of the two industries in the United States, coupled with a shift of workers toward the other in Germany, could result in increased total production of both goods in the two countries.

(d) Suppose that when trade opens, the international price ratio settles at the former German price ratio. Using diagrams, show what happens to production, consumption, exports, and imports (i) in the United States and (ii) in Germany.
Answer #1

a) A worker in US can produce 1/20 = 0.05 cars and a worker in Germany can produce 1/10=0.1 cars. Also, a worker in US can produce 1000/25=40 chickens and a worker in German can produce 1000/20=50 chickens.

Now, opportunity cost of producing Cars in US is 40/0.05=800 and that in Germany is 50/0.1 = 500. As opportunity cost of producing cars is lower in Germany, it has comparative advantage in production of cars. Hence, workers in Germany must specialize in car production.

Similarly, opportunity cost of producing Chickens in US is 0.05/40=0.00125 and that in Germany is 0.1/50=0.002. As opportunity cost of producing chickens is lower in US, it has comparative advantage in production of chickens. Hence, workers in US must specialize in chicken production.

b) Now, when free trade takes place at German price , i.e, autarky price of Germany , (Pcars/Pchickens)Ger = 500 chickens per car.

Then, in US, for 40 chickens produced by a worker, it will get 1/500*40 = 0.08 cars. Hence, gains from trade is higher for US.

Similarly, in Germany, for 0.1 cars produced by a worker, it will get 500*0.1 = 50 chickens. Hence, trade conditions remains unchanged for Germany.

In the diagram below we have shown the CPF and PPF for both countries.

Cars Cars 0.1 0.08 CPF PPF for Germany CPF for US 0.05 PPF for US 50 40 Chickens Chickens